I was given a right triangle. One side of the triangle is x, the hypotenus is x+5 and the base of the triangle is 15. There is no degrees, and it says to find the value of X.
Answer by ankor@dixie-net.com(22740)

a right triangle. One side of the triangle is x, the hypotenuse is x+5 and the base of the triangle is 15. find the value of X.
:
Using Pythagoras a^2 + b^2 = c^2
In this problem:
a = x
b = 15
c = (x+5)
:
x^2 + 15^2 = (x+5)^2
x^2 + 225 = x^2 + 10x + 25
x^2 - x^2 + 225 - 25 = 10x
200 = 10x
x = 200/10
x = 20
:
:
See if this works
20^2 + 15^2 = 25^2